Find four consecutive even integers such that if the sum of the first and third is multiplied by 4 the result is 10 more than 7 times the fourth.
Answer by ankor@dixie-net.com(22740)

Find four consecutive even integers
x, (x+2), (x+4), (x+6)
:
such that if the sum of the first and third is multiplied by 4 the result is 10 more than 7 times the fourth.
4[x + (x+4)] = 7(x+6) + 10
:
4(2x+4) = 7x + 42 + 10
:
8x + 16 = 7x + 52
:
8x - 7x = 52 - 16
:
x = 36 is he 1st integer
:
36, 38, 40, 42 the four consecutive integers
:
:
Check this in the statement:
"sum of the first and third is multiplied by 4 the result is 10 more than 7 times the fourth"
4(36 + 40) = 7(42) + 10
304 = 294 + 10